# Impact of space-time curvature coupling on the vacuum energy induced by a magnetic topological defect in flat space-time of arbitrary dimension

### v1

@@ Abstract
We have investigated vacuum polarization of a quantized charged massive scalar field in the presence of a magnetic topological defect, modeled as an impenetrable tube of finite thickness carrying magnetic flux. At the tube’s surface, we imposed a generalized Robin boundary condition. Our analysis demonstrates that, in flat space-time, the total induced vacuum energy is independent of the coupling $\xi$ of the scalar field’s interaction with the space-time curvature only in the special cases of Dirichlet and Neumann boundary conditions. For general Robin boundary conditions, however, the total induced vacuum energy depends on the coupling $\xi$ in a flat space-time and exhibits a nontrivial dependence on the parameter of the Robin boundary condition. We investigated the dependence of this effect not only on Robin’s boundary condition parameter, but also on the tube thickness and the space-time dimensionality. We conclude that careful measurements of vacuum polarization effects in flat space-time may, in principle, provide an independent way to probe the $\xi$ coupling.
@@ 1 Introduction
In this paper, we study vacuum polarization by a magnetic topological defect, modeled as an infinitely long cylindrical tube that is impenetrable to a matter field and contains a magnetic flux inside. In such a configuration, the Aharonov-Bohm effect [6] plays a crucial role, as the enclosed magnetic field can influence the quantum matter field located outside the tube. Within the framework of second-quantized field theory, such a topological defect gives rise to an induced vacuum energy, vacuum current and the corresponding magnetic flux in the exterior region (see, e.g., the pioneering studies carried out in the singular magnetic vortex approximation [7, 8, 9, 10]). The imposed boundary condition at the tube’s surface has a significant impact on the behavior of the matter field outside. Since the observed quantum effects originate from both the boundary condition and the internal magnetic flux, this phenomenon is commonly referred to as the Casimir–Bohm–Aharonov effect [11].
The phenomenon of vacuum polarization near the linear magnetic topological defects has a wide range of physical applications. These topological defects can be derived from phase transitions with spontaneous gauge symmetry breaking in the early Universe or condensed matter physics. The impenetrable magnetic tube model can be applied to describe magnetic cosmic strings, the formation of which was possible in the early Universe [12, 13, 14, 15]. In condensed matter physics, the impenetrable magnetic tube can be considered as a model of Abrikosov-Nielsen-Olesen vortex in type-II superconductors [16, 17] or as disclinations in nanoconical structures of two-dimensional materials like graphene, see, e.g., [18, 19, 20, 21, 22, 23].
@@
In the case of $(3+1)$-dimensional space-time $\xi_{c}=1/6$. It was shown in [33] that the value $\xi=1/4$ also plays an important role, which we will discuss further.
In this paper, we will investigate the vacuum polarization of a charged massive scalar matter field $\psi$ in the flat space-time of arbitrary dimension in the background of a magnetic tube of finite transverse size $r_{0}$, which is impenetrable to the matter field. We will be interested in the dependence of the induced vacuum energy on the coupling to the space-time curvature. On the surface of the tube, we consider generalized Robin-type boundary conditions
@@
where $\theta$ is a parameter of the boundary condition, $\pi/2\leq\theta<\pi/2$. The special cases of
$\theta=0$ and $\theta=-\pi/2$ correspond to Dirichlet and Neumann boundary conditions, respectively. For these special cases, the induced vacuum energy has been studied in [34, 35, 36] (with the Dirichlet boundary condition) and in [37] (with the Neumann boundary condition).
@@
We will use the Lagrangian (1) for a complex scalar massive field $\psi$ in $(d+1)$-dimensional space-time. In the case of a static background [27, 39, 40], the operator of the quantized charged scalar field has the form
@@
where $a^{\dagger}_{\lambda}$ and $a_{\lambda}$ ($b^{\dagger}_{\lambda}$ and $b_{\lambda}$) are the scalar particle (antiparticle) creation and annihilation operators; $\lambda$ is the set of parameters (quantum numbers) specifying the state; $E_{\lambda}=E_{-\lambda}>0$ is the energy of the state; symbol $\sum\hskip-10.00002pt\int\limits_{\lambda}$ denotes summation over discrete and integration (with a certain measure) over continuous values of $\lambda$; wave functions $\psi_{\lambda}(\textbf{x})$ are the solutions to the stationary equation of motion,
$\left\{-{\mbox{\boldmath$\nabla$}}^{2}+(m^{2}+\xi R)\right\}\psi_{\lambda}(\textbf{x})=E^{2}_{\lambda}\psi(\textbf{x}),$ (7)
@@
$\varepsilon=\langle\rm vac|\left[\partial_{0}\Psi^{\dagger}\partial_{0}\Psi+\partial_{0}\Psi\partial_{0}\Psi^{\dagger}+(1/4-\xi)\mbox{\boldmath$\nabla$}^{2}(\Psi^{\dagger}\Psi+\Psi\Psi^{\dagger})\right]-\xi R^{00}(\Psi^{\dagger}\Psi+\Psi\Psi^{\dagger})|\rm vac\rangle.$ (8)
We now focus on the case of flat space-time, where $R^{\mu\nu}=0$, $R=0$. In this case, the vacuum energy density still depends on the coupling $\xi$ and can be written as
$\varepsilon=\varepsilon_{can}+(1/4-\xi)\varepsilon_{\xi}=\sum\hskip-13.99995pt\int\limits_{\lambda}E_{\lambda}\psi^{*}_{\lambda}(\textbf{x})\,\psi_{\lambda}(\textbf{x})+(1/4-\xi)\mbox{\boldmath$\nabla$}^{2}\sum\hskip-13.99995pt\int\limits_{\lambda}E^{-1}_{\lambda}\psi^{*}_{\lambda}(\textbf{x})\,\psi_{\lambda}(\textbf{x}),$ (9)
@@
We consider a static background described by a cylindrically symmetric tube of finite transverse size carrying a magnetic flux. The coordinate system is chosen such that the tube is oriented along the $z$-axis. In $(d+1)$-dimensional space-time, this configuration can be naturally generalized to the (d2)(d-2)-tube by adding an additional d3d-3 longitudinal directions. The covariant derivative is defined as $\nabla_{0}=\partial_{0}$, $\mbox{\boldmath$\nabla$}=\mbox{\boldmath$\partial$}-{\rm i}\tilde{e}\,{\bf V}$ where $\tilde{e}$ is the coupling of dimension $m^{(3-d)/2}$, and the vector potential has only one nonvanishing component,
@@
$0<k<\infty$, $\infty<p^{j}<\infty$ for $j=1,\dots,d{-}2$; and $n\in\mathbb{Z}$, where $\mathbb{Z}$ denotes the set of integer numbers. The functions $J_{\rho}(u)$ and $Y_{\rho}(u)$ are the Bessel functions of the first and second kinds of order $\rho$, respectively. A prime on a function denotes differentiation with respect to its argument. The solutions given by Eq. (11) satisfy the orthonormalization condition
@@
Unfortunately, the Exp.(9) exhibits ultraviolet divergences and requires renormalization, see, e.g., [3]. In our case, where the magnetic field is excluded from the region accessible to the matter field, the renormalization procedure reduces to a single subtraction: the removal of the contribution corresponding to the absence of the magnetic flux, see [41].
@@
As a result of the infinite range of summation, the function $S$ and the induced quantities depend only on the fractional magnetic flux inside the tube
@@ 3 Induced vacuum energy in $2+1$ space-time
$\varepsilon_{ren}=\frac{1}{2\pi}\int\limits_{0}^{\infty}dk\,k\left(\sqrt{k^{2}+m^{2}}+\frac{1/4-\xi}{\sqrt{k^{2}+m^{2}}}\triangle_{r}\right)G(\theta,kr,kr_{0},\Phi),$ (20)
where $\triangle_{r}=\partial^{2}_{r}+r^{-1}\partial_{r}$ is the transverse radial part of the Laplacian.
@@
Figure 1: The total induced dimensionless vacuum energy $E_{\xi}^{2+1}/m$ (26) in $(2+1)$-dimensional space-time as the function of parameter $\theta$ of Robin boundary conditions on the edge of the magnetic impenetrable tube is presented for different thicknesses of the tube. For convenience, the above functions are multiplied by the coefficient $c$.
@@ 4 Induced vacuum energy in higher-dimensional spaces
As shown in the previous sections, the induced vacuum energy density depends solely on the distance from the string within the plane transverse to the tube. Accordingly, in higher-dimensional space-times, the total induced energy becomes divergent due to integration over directions perpendicular to this transverse plane. Therefore, it is meaningful to consider the induced energy only within the plane transverse to the tube.
@@
$E^{(d+1)}_{can}=(2\pi)^{2-d}\int\limits_{r_{0}}^{\infty}dr\,r\int\limits_{-\infty}^{\infty}d^{d-2}{\bf p}\int\limits_{0}^{\infty}dk\,k\sqrt{{\bf p}^{2}+k^{2}+m^{2}}\,G(\theta,kr,kr_{0},F)=\\ =\frac{2}{r_{0}^{d-1}}\frac{(4\pi)^{1-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}\int\limits_{x_{0}}^{\infty}dv\,v^{2}\left(v^{2}-x_{0}^{2}\right)^{\frac{d-4}{2}}\,\mathcal{D}_{can}\left(\theta,v\right),$ (27)
@@
$E^{(2+1)}_{can}=m\mathcal{D}_{can}(\theta,mr_{0}),$ (28)
$\mathcal{D}_{can}(\theta,y)=\int\limits_{y}^{\infty}\frac{dx}{x^{2}}\int\limits_{0}^{\infty}dz\,z\sqrt{z^{2}+x^{2}}\,G(\theta,z,zy/x,F).$ (29)
@@
$E^{(d+1)}_{\xi}=(2\pi)^{2-d}\int\limits_{r_{0}}^{\infty}dr\,r\int\limits_{-\infty}^{\infty}d^{d-2}{\bf p}\int\limits_{0}^{\infty}dk\,k\,\frac{\Delta_{r}G(\theta,kr,kr_{0},F)}{\sqrt{{\bf p}^{2}+k^{2}+m^{2}}}=\\ =\frac{2}{r_{0}^{d-1}}\frac{(4\pi)^{1-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}\int\limits_{x_{0}}^{\infty}dv\,v^{2}\left(v^{2}-x_{0}^{2}\right)^{\frac{d-4}{2}}\,\mathcal{D}_{\xi}(\theta,v),$ (30)
@@
$E^{(2+1)}_{\xi}=m\mathcal{D}_{\xi}(\theta,mr_{0}),$ (31)
$\mathcal{D}_{\xi}(\theta,y)=\int\limits_{y}^{\infty}dx\,x\Delta_{x}\int\limits_{0}^{\infty}dz\,z\frac{G(\theta,z,zy/x,F)}{x\sqrt{z^{2}+x^{2}}}.$ (32)
@@
$E^{(d+1)}(mr_{0})=E^{(d+1)}_{can}(mr_{0})+(1/4-\xi)E^{(d+1)}_{\xi}(mr_{0})=\\ =\frac{2}{r_{0}^{d-1}}\frac{(4\pi)^{1-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}\int\limits_{x_{0}}^{\infty}dv\,v^{2}\left(v^{2}-x_{0}^{2}\right)^{\frac{d-4}{2}}\left[\mathcal{D}_{can}\left(\theta,v\right)+(1/4-\xi)\mathcal{D}_{\xi}\left(\theta,v\right)\right]$ (33)
@@
The results of our computations for the function $E^{(d+1)}_{\xi}$ (30) of primary interest to us, for the specific cases $d=3$ and $d=4$, are presented in Fig.2. One can see that the induced vacuum dimensionless energy in $d+1$ space-time $E^{(d+1)}_{\xi}/m^{d-1}$ is rapidly increasing with increasing space dimension $d$ for sufficiently small transverse sizes of the topological defect $mr_{0}$. The main question is: how large are the values of the $E^{(d+1)}_{\xi}/m^{d-1}$ functions? To answer this, let’s compare them with the values of the $E^{(d+1)}_{can}/m^{d-1}$ functions.
Figure 2: The induced dimensionless vacuum energy $E_{\xi}^{d+1}/m^{d-1}$ (30) within the plane transverse to the tube in $(d+1)$-dimensional space-time as the function of parameter $\theta$ of Robin boundary conditions on the edge of the magnetic impenetrable tube is presented for different thicknesses of the tube: a) $d=3$, b) $d=4$. For convenience, the above functions are multiplied by the coefficient $c$.
@@
Using the results of [36, 37], we reconstruct the dependence of the induced dimensionless vacuum energy $E^{(d+1)}_{can}/m^{d-1}$ on the tube thickness for the Dirichlet and Neumann cases, see Fig.3, and compare it with the maximal value of $E^{(d+1)}_{\xi}/m^{d-1}$ attained at some $\theta_{\mathrm{max}}$ in the range $\pi/2<\theta<0$, see Table 1. The value of $E^{(d+1)}_{can}/m^{d-1}$ at $\theta_{\mathrm{max}}$ lies between those for the Dirichlet and Neumann boundary conditions. As seen from Table 1, for certain values of the boundary condition parameter $\pi/2<\theta<0$, the quantity $E^{(d+1)}_{\xi}/m^{d-1}$ is of the same order of magnitude, or even larger, than $E^{(d+1)}_{can}/m^{d-1}$ for the same $\theta$. It has been shown that this statement is true for cases of different dimensions of space-time.
Figure 3: The induced dimensionless vacuum energy $E_{\xi}^{d+1}/m^{d-1}$ within the plane transverse to the tube in $(d+1)$-dimensional space-time as the function of the thicknesses of the magnetic impenetrable tube $mr_{0}$ for the cases of a) Dirichlet and b) Neumann boundary conditions.

### v2

@@ Abstract
We have investigated vacuum polarization of a quantized charged massive scalar field in the presence of a magnetic topological defect, modeled as an impenetrable tube of finite thickness carrying magnetic flux. At the tube’s surface, we imposed a general Robin boundary condition. Our analysis demonstrates that, in flat space-time, the total induced vacuum energy is independent of the coupling $\xi$ of the scalar field’s interaction with the space-time curvature only in the special cases of Dirichlet and Neumann boundary conditions. For general Robin boundary conditions, however, the total induced vacuum energy depends on the coupling $\xi$ in a flat space-time and exhibits a nontrivial dependence on the parameter of the Robin boundary condition. We investigated the dependence of this effect not only on Robin’s boundary condition parameter, but also on the tube thickness and the space-time dimensionality. We conclude that careful measurements of vacuum polarization effects in flat space-time may, in principle, provide an independent way to probe the $\xi$ coupling.
@@ 1 Introduction
In this paper, we study vacuum polarization by a magnetic topological defect, modeled as an infinitely long cylindrical tube that is impenetrable to a matter field and contains a magnetic flux inside. In such a configuration, the Aharonov-Bohm effect [2] plays a crucial role, as the enclosed magnetic field can influence the quantum matter field located outside the tube. By vacuum polarization we mean the modification of the quantum vacuum caused by an external field or boundary, which alters the spectrum of virtual particles and leads to changes in observable quantities. Within the framework of second-quantized field theory, such a topological defect gives rise to an induced vacuum energy, vacuum current and the corresponding magnetic flux in the exterior region (see, e.g., the pioneering studies carried out in the singular magnetic vortex approximation [39, 23, 16, 36]). The imposed boundary condition at the tube’s surface has a significant impact on the behavior of the matter field outside. Since the observed quantum effects originate from both the boundary condition and the internal magnetic flux, this phenomenon is commonly referred to as the Casimir–Bohm–Aharonov effect [40].
The phenomenon of vacuum polarization near the linear magnetic topological defects has a wide range of physical applications. These topological defects can be derived from phase transitions with spontaneous gauge symmetry breaking in the early Universe or condensed matter physics. The impenetrable magnetic tube model can be applied to describe magnetic cosmic strings, the formation of which was possible in the early Universe [29, 48, 47, 26]. In condensed matter physics, the impenetrable magnetic tube can be considered as a model of Abrikosov-Nielsen-Olesen vortex in type-II superconductors [1, 34] or as disclinations in nanoconical structures of two-dimensional materials like graphene, see, e.g., [31, 25, 41, 33, 44, 45, 4].
@@
In the case of $(3+1)$-dimensional space-time $\xi_{c}=1/6$. It was shown in [43] that the value $\xi=1/4$ also plays an important technical role, which we will discuss further.
In this paper, we will investigate the vacuum polarization of a charged massive scalar matter field $\psi$ in the flat space-time of arbitrary dimension in the background of a magnetic tube of finite transverse size $r_{0}$, which is impenetrable to the matter field. We will be interested in the dependence of the induced vacuum energy on the coupling to the space-time curvature. On the surface of the tube, we consider general Robin-type boundary conditions
@@
where $r_{0}$ is the radius of the tube and $\theta$ is a parameter of the boundary condition, $\pi/2\leq\theta<\pi/2$. The special cases of $\theta=0$ and $\theta=-\pi/2$ correspond to Dirichlet and Neumann boundary conditions, respectively. For these special cases, the induced vacuum energy has been studied in [19, 21, 22] (with the Dirichlet boundary condition) and in [18] (with the Neumann boundary condition).
@@
We will use the Lagrangian (1) for a complex scalar massive field $\psi$ in $(d+1)$-dimensional space-time. In the case of a static background [5, 24, 35], the operator of the quantized charged scalar field, whose classical version appears in Lagrangian (1), has the form
@@
where $a^{\dagger}_{\lambda}$ and $a_{\lambda}$ ($b^{\dagger}_{\lambda}$ and $b_{\lambda}$) are the scalar particle (antiparticle) creation and annihilation operators; $\lambda$ is the set of parameters (quantum numbers) specifying the state; $E_{\lambda}=E_{-\lambda}>0$ is the energy of the state; symbol $\sum\hskip-10.00002pt\int\limits_{\lambda}$ denotes summation over discrete and integration (with a certain measure) over continuous values of $\lambda$; wave functions $\psi_{\lambda}(\textbf{x})$ are the solutions to the stationary equation of motion that follow from Lagrangian (1),
$\left\{-{\mbox{$\nabla$}}^{2}+(m^{2}+\xi R)\right\}\psi_{\lambda}(\textbf{x})=E^{2}_{\lambda}\psi(\textbf{x}),$ (7)
@@
$\varepsilon=\langle\rm vac|\left[\partial_{0}\Psi^{\dagger}\partial_{0}\Psi+\partial_{0}\Psi\partial_{0}\Psi^{\dagger}+(1/4-\xi)\mbox{$\nabla$}^{2}(\Psi^{\dagger}\Psi+\Psi\Psi^{\dagger})\right]-\xi R^{00}(\Psi^{\dagger}\Psi+\Psi\Psi^{\dagger})|\rm vac\rangle,$ (8)
where the vacuum state $|\mathrm{vac}\rangle$ is defined in the usual way by the conditions $a_{\lambda}|\rm vac\rangle=b_{\lambda}|\rm vac\rangle=0$. We now focus on the case of flat space-time, where $R^{\mu\nu}=0$, $R=0$. In this case, the vacuum energy density still depends on the coupling $\xi$ and can be written as
$\varepsilon=\varepsilon_{can}+(1/4-\xi)\varepsilon_{\xi}=\sum\hskip-13.99995pt\int\limits_{\lambda}E_{\lambda}\psi^{*}_{\lambda}(\textbf{x})\,\psi_{\lambda}(\textbf{x})+(1/4-\xi)\mbox{$\nabla$}^{2}\sum\hskip-13.99995pt\int\limits_{\lambda}E^{-1}_{\lambda}\psi^{*}_{\lambda}(\textbf{x})\,\psi_{\lambda}(\textbf{x}),$ (9)
@@
[…]
We consider a static background described by a cylindrically symmetric tube of finite transverse size carrying a magnetic flux. The coordinate system is chosen such that the tube is oriented along the $z$-axis. In $(d+1)$-dimensional space-time, this configuration can be naturally generalized to the (d1)(d-1)-tube by adding an additional d3d-3 longitudinal directions. Thus, in $2+1$ space-time, we have a ring lying in the plane perpendicular to the $z$-axis; in $3+1$ space-time, we obtain a tube extending along the $z$-axis; and in $4+1$ space-time, this generalises to a hypersurface consisting of all points that satisfy $x^{2}+y^{2}=r_{0}^{2}$ at arbitrary values of $z$ and additional $x^{1}$ coordinates. The covariant derivative is defined as $\nabla_{0}=\partial_{0}$, $\mbox{$\nabla$}=\mbox{$\partial$}-{\rm i}\tilde{e}\,{\bf V}$ where $\tilde{e}$ is the coupling of dimension $m^{(3-d)/2}$, and the vector potential has only one nonvanishing component,
@@
$0<k<\infty$, $\bf{p}=(p^{z},p^{1},...,p^{d-3})$, $\bf{x}_{d-2}=(z,x^{1},...,x^{d-3})$, $\infty<p^{z}<\infty$, $\infty<p^{j}<\infty$ for $j=1,\dots,d{-}3$; and $n\in\mathbb{Z}$, where $\mathbb{Z}$ denotes the set of integer numbers. The functions $J_{\rho}(u)$ and $Y_{\rho}(u)$ are the Bessel functions of the first and second kinds of order $\rho$, respectively. A prime on a function denotes differentiation with respect to its argument. The solutions given by Eq. (11) satisfy the orthonormalization condition
@@
Unfortunately, the Exp.(9) exhibits ultraviolet divergences and requires renormalization, see, e.g., [32]. In our setting, where the magnetic field is excluded from the region accessible to the matter field, the renormalization procedure reduces to a single subtraction: removing the contribution corresponding to the absence of magnetic flux, see [3]. Consequently, the induced vacuum energy is defined as the change in vacuum energy between an impenetrable tube carrying a magnetic flux and the same impenetrable tube without flux. Of course, both quantities also contain a surface divergence arising from the presence of the boundary (the tube surface) - the Casimir surface energy. This contribution is determined solely by the geometry/topology of the boundary and by the boundary conditions imposed on the fields [32, 6, 7, 46]; it is independent of the presence of the magnetic flux inside the tube. Consequently, this divergence cancels upon subtraction and does not require the introduction of additional counterterms.
@@
As can be seen from (16), the value $\xi=1/4$ is technically distinguished: it reduces the energy density to its canonical expression. The factor $(1/4-\xi)$ separates the terms in the renormalized energy according to their convergence behavior at large integration momenta. Furthermore, note that the induced energy density in $(d+1)$-dimensional space-time does not depend on additional spatial coordinates.
As a result of the infinite range of summation, the function $S$ and consequently the function $G$, as well as the induced quantities, depend only on the fractional magnetic flux inside the tube
@@ 3 Induced vacuum energy in $2+1$ space-time
$\varepsilon_{ren}=\frac{1}{2\pi}\int\limits_{0}^{\infty}dk\,k\left(\sqrt{k^{2}+m^{2}}+\frac{1/4-\xi}{\sqrt{k^{2}+m^{2}}}\triangle_{r}\right)G(\theta,kr,kr_{0},{F}),$ (20)
where $\triangle_{r}=\partial^{2}_{r}+r^{-1}\partial_{r}$ is the transverse radial part of the Laplacian. We have also redefined the function $G$, and in what follows, we will consider it to be expressed in terms of $F$ as specified in (19).
@@
Figure 1: The total induced dimensionless vacuum energy $E_{\xi}^{2+1}/m$ (26) in $(2+1)$-dimensional space-time as the function of parameter $\theta$ of Robin boundary conditions on the edge of the magnetic impenetrable tube is presented for different thicknesses of the tube. For convenience, the above functions are multiplied by the coefficient $c$. For example, the value of the function presented in the figure for $mr_{0}=1/10$ is multiplied by a factor of $c=200$.
@@ 4 Induced vacuum energy in higher-dimensional spaces
As shown in the previous sections, the induced vacuum energy density depends solely on the distance from the string within the plane transverse to the tube. Accordingly, in higher-dimensional space-times, the total induced energy becomes divergent due to integration of the induced energy density (16) over coordinates perpendicular to this transverse plane. Therefore, it is meaningful to consider the induced energy only within the plane transverse to the tube.
@@
$E^{(d+1)}_{can}=(2\pi)^{2-d}\int\limits_{r_{0}}^{\infty}dr\,r\int\limits_{-\infty}^{\infty}d^{d-2}{\bf p}\int\limits_{0}^{\infty}dk\,k\sqrt{{\bf p}^{2}+k^{2}+m^{2}}\,G(\theta,kr,kr_{0},F)=\\ =\frac{2}{r_{0}^{d-1}}\frac{(4\pi)^{1-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}\int\limits_{x_{0}}^{\infty}dv\,v^{2}\left(v^{2}-x_{0}^{2}\right)^{\frac{d-4}{2}}\,\mathcal{D}_{can}\left(F,\theta,v\right),$ (27)
@@
$E^{(2+1)}_{can}=m\mathcal{D}_{can}(F,\theta,mr_{0}),$ (28)
$\mathcal{D}_{can}(F,\theta,y)=\int\limits_{y}^{\infty}\frac{dx}{x^{2}}\int\limits_{0}^{\infty}dz\,z\sqrt{z^{2}+x^{2}}\,G(\theta,z,zy/x,F).$ (29)
Figure 2: The induced dimensionless vacuum energy $E_{\xi}^{d+1}/m^{d-1}$ (30) within the plane transverse to the tube in $(d+1)$-dimensional space-time as the function of parameter $\theta$ of Robin boundary conditions on the edge of the magnetic impenetrable tube is presented for different thicknesses of the tube: a) $d=3$, b) $d=4$. For convenience, the above functions are multiplied by the coefficient $c$. For example, the value of the function presented in the figure a) for $mr_{0}=10^{-1}$ is multiplied by a factor of $c=8000$.
@@
$E^{(d+1)}_{\xi}=(2\pi)^{2-d}\int\limits_{r_{0}}^{\infty}dr\,r\int\limits_{-\infty}^{\infty}d^{d-2}{\bf p}\int\limits_{0}^{\infty}dk\,k\,\frac{\Delta_{r}G(\theta,kr,kr_{0},F)}{\sqrt{{\bf p}^{2}+k^{2}+m^{2}}}=\\ =\frac{2}{r_{0}^{d-1}}\frac{(4\pi)^{1-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}\int\limits_{x_{0}}^{\infty}dv\,v^{2}\left(v^{2}-x_{0}^{2}\right)^{\frac{d-4}{2}}\,\mathcal{D}_{\xi}(F,\theta,v),$ (30)
@@
$E^{(2+1)}_{\xi}=m\mathcal{D}_{\xi}(F,\theta,mr_{0}),$ (31)
$\mathcal{D}_{\xi}(F,\theta,y)=\int\limits_{y}^{\infty}dx\,x\Delta_{x}\int\limits_{0}^{\infty}dz\,z\frac{G(\theta,z,zy/x,F)}{x\sqrt{z^{2}+x^{2}}},$ (32)
where $\triangle_{r}=\partial^{2}_{x}+x^{-1}\partial_{x}$.
@@
$E^{(d+1)}=E^{(d+1)}_{can}+(1/4-\xi)E^{(d+1)}_{\xi}=\\ =\frac{2}{r_{0}^{d-1}}\frac{(4\pi)^{1-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}\int\limits_{x_{0}}^{\infty}dv\,v^{2}\left(v^{2}-x_{0}^{2}\right)^{\frac{d-4}{2}}\left[\mathcal{D}_{can}\left(F,\theta,v\right)+(1/4-\xi)\mathcal{D}_{\xi}\left(F,\theta,v\right)\right]$ (33)
@@
The results of our computations for the function $E^{(d+1)}_{\xi}$ (30) are shown in Fig.2 for the representative cases $d=3$ and $d=4$ , and half-integer magnetic flux $F=1/2$. One can see that the total energy (33), in addition to the canonical contribution, contains a term $(1/4-\xi)\,E^{(d+1)}_{\xi}$, and the quantity $E^{(d+1)}_{\xi}$ does not vanish in spaces of arbitrary dimension. The value of $E^{(d+1)}_{\xi}$ is determined by the parameter $\theta$ of the general Robin boundary condition and exhibits a similar dependence on this parameter across different space–time dimensions. Specifically, the induced energy $E^{(d+1)}_{\xi}$ vanishes at the endpoints of the interval $\pi/2\leq\theta\leq 0$, while inside this interval it is a smooth function with a single maximum. One can also see that the induced vacuum dimensionless energy in $d+1$ space-time $E^{(d+1)}_{\xi}/m^{d-1}$ is rapidly increasing with increasing space dimension $d$ for sufficiently small transverse sizes of the topological defect $mr_{0}$ and rapidly increasing as $mr_{0}$ decreases. The main question is: how large are the values of the $E^{(d+1)}_{\xi}/m^{d-1}$ functions? To answer this, let’s compare them with the values of the $E^{(d+1)}_{can}/m^{d-1}$ functions.
@@
Using the results of [22, 18], we reconstruct the dependence of the induced dimensionless vacuum energy $E^{(d+1)}_{can}/m^{d-1}$ on the tube thickness for the Dirichlet and Neumann cases, see Fig.3, and compare it with the maximal value of $E^{(d+1)}_{\xi}/m^{d-1}$ attained at some $\theta_{\mathrm{max}}$ in the range $\pi/2<\theta<0$, see Table 1. The value of $E^{(d+1)}_{\xi}/m^{d-1}$ at $\theta_{\mathrm{max}}$ lies between those for the Dirichlet and Neumann boundary conditions. As seen from Table 1, for certain values of the boundary condition parameter $\pi/2<\theta<0$, the quantity $E^{(d+1)}_{\xi}/m^{d-1}$ is of the same order of magnitude, or even larger, than $E^{(d+1)}_{can}/m^{d-1}$ for the same $\theta$. It has been shown that this statement is true for cases of different dimensions of space-time.
Figure 3: The induced dimensionless vacuum energy $E_{can}^{d+1}/m^{d-1}$ within the plane transverse to the tube in $(d+1)$-dimensional space-time as the function of the thicknesses of the magnetic impenetrable tube $mr_{0}$ for the cases of a) Dirichlet and b) Neumann boundary conditions.
We would also like to note that Fig.3 shows the dependence of the canonical contribution to the induced energy, $E^{(d+1)}_{can}$, on the thickness of the magnetic tube for a fixed boundary-condition parameter. One can see that $E^{(d+1)}_{can}$ increases rapidly as $mr_{0}$ decreases, and the same behavior is exhibited by the $E^{(d+1)}_{\xi}$ contribution in Fig.2. However, Fig.3 additionally allows one to see that $E^{(d+1)}_{can}$ increases with increasing space dimension $d$ only for small tube radius, $mr_{0}<mr_{c}$, and decreases for large tube radius, $mr_{0}>mr_{c}$. For the cases $d=2,3,4$, we have $mr_{c}\approx 0.05$, and it grows slowly as the spatial dimension increases. The same behavior can be observed from a detailed analysis of Fig.2 and Table 1 for the $E^{(d+1)}_{\xi}$ contribution to the induced energy.